OpenStudy (anonymous):

I need an answer checked! I'll fan and medal!! Simplify 25/55x A: 5/55x B: 1/2x C: x/2 D: 5x/12 The answer I got is 5/11x, Is that correct?

OpenStudy (anonymous):

A is 5/11x oops

OpenStudy (gabebae):

It's correct

OpenStudy (alexandervonhumboldt2):

yes you are correct

OpenStudy (anonymous):

Thank you!

OpenStudy (alexandervonhumboldt2):

\[\frac{ 25 }{ 55 }=\frac{ 5*5 }{ 5*11 }=5/11\]